OpenStudy (anonymous):

Find the numerical value if x=1 and y=2.

OpenStudy (anonymous):

\[\left(\begin{matrix}2x+y^2 \\ 2x+2y\end{matrix}\right)\]

OpenStudy (anonymous):

Plug the x values and y values into the equation

OpenStudy (anonymous):

oh that is it?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok can you help me solve this one then... 6(x+4)=8(x-2)